Futuristic was a newly setup firm in 2019. It is in the business of providing artwork. However, being in the business of standard art objects (non-essential items), it did not do well once it was hit by COVID-19. It has a selling price of ₹2,500 per piece and a variable cost of ₹1,000 per piece. It incurs an annual fixed cost of ₹30,00,000.

a. The manager of the business wishes to know-  the number of art pieces they must sell to be at the break-even point  the contribution margin


1
Expert's answer
2021-10-03T18:17:18-0400

The break-even point is:

"BEP = 3,000,000\/(2,500 - 1,000) = 2,000" units.

The contribution margin is:

"CM = (2,500 - 1,000)\/2,500 = 0.6."